A tannery with a wastewater flow of 0.011 m^3/s an
Answer #1

Solution.

The ultimate BOD of tannery wastewater is

L_0 = \frac{590}{1-e^{-0.115\cdot 5}} = 1349.2 \ \frac{mg}{L};

The ultimate BOD of a creek is

L_0 = \frac{0.6}{1-e^{-3.7\cdot 5}} = 0.6 \ \frac{mg}{L};

The initial ultimate BOD can be calculated using an equation

L_a = \frac{Q_wL_w+Q_rL_r}{Q_w+Q_r};

L_a = \frac{0.011\cdot 1349.2+1.7\cdot 0.6}{0.011+1.7} = 9.27 \ \frac{mg}{L}.